OpenStudy (anonymous):

olve the following system of equations by the substitution method. 10x + 10y = 1 x = y - 3 What is the value of y? -29/20 31/20 -3/2

OpenStudy (anonymous):

See, x = y - 3 Put this value in First Equation: 10(y - 3) + 10y = 1 10y - 30 + 10y = 1 20y = 1 + 30 20y = 31 y = 31/20 is the answer...

OpenStudy (anonymous):

thx i had got confused y it was 31/20 but its because u had to divid 31by 20